A player kicks a soccer ball towards the goalkeeper, but at an angle of 36degrees to the horizontal with an initial speed speed of 14.7m/s. The goalkeeper stands 26.0m from the kicker. Where will the ball land relative to the goalkeeper?
1
Expert's answer
2019-11-11T15:52:32-0500

The range of projectile:


"R=\frac{v^2}{g}\sin{2(36)}"

"R=\frac{14.7^2}{9.8}\sin{72}=21.0\ m""s=d-R=26-21=5.0\ m"

The ball lands 5.0 m before the goalkeeper.